OpenStudy (anonymous):

a particle moves along a line so that, at time t, its position is s(t)=8sin2t. for what values of t does the particle change direction?

OpenStudy (anonymous):

where \(\cos(2t)=0\) because that is where it goes from increasng to decreasing

OpenStudy (anonymous):

okay which is pi/4

OpenStudy (anonymous):

but how would i figure out the particles max velcoity? (Next part of the question)

OpenStudy (anonymous):

that is one point, yes

OpenStudy (anonymous):

velocity is given by \(16\cos(2t)\) as it is the derivative of the position. differentiate again to get the second derivative, that will tell you where the first derivative has a max

OpenStudy (anonymous):

which is -32sin2t?

OpenStudy (anonymous):

actually if it just says "what is the particles max velocity" you can reason that since \(-1\leq \cos(x)\leq 1\) you must have the maximimum velocity is 16

OpenStudy (anonymous):

since that is the biggest \(16\cos(2t)\) can be

OpenStudy (anonymous):

and how did u determine it is 16?

OpenStudy (anonymous):

yes your second derivative is right

OpenStudy (anonymous):

because the largest cosine can be is 1 so the largest 16cosine can be is 16

OpenStudy (anonymous):

how would i show my work for that?

OpenStudy (anonymous):

i would write that since \(-1\leq \cos(x)\leq 1\) you get \(-16\leq16\cos(2t)\leq 16\)

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

how would i determine the max distance from the origin for the function, s=4sin4t